### v1

# Basis-Spline Assisted Coded Computing: Strategies and Error Bounds

Abstract
Coded computing has become a key framework for reliable distributed computation over decentralized networks, effectively mitigating the impact of stragglers. Although there exists a wide range of coded computing methods to handle both polynomial and non-polynomial functions, computing methods for the latter class have received traction due its inherent challenges in reconstructing non-polynomial functions using a finite number of evaluations. Among them, the state-of-the-art method is Berrut Approximated coded computing, wherein Berrut interpolants, are used for approximating the non-polynomial function. However, since Berrut interpolants have global support characteristics, such methods are known to offer degraded accuracy when the number of stragglers is large. To address this challenge, we propose a coded computing framework based on cubic B-spline interpolation. In our approach, server-side function evaluations are reconstructed at the master node using B-splines, exploiting their local support and smoothness properties to enhance stability and accuracy. We provide a systematic methodology for integrating B-spline interpolation into coded computing and derive theoretical bounds on approximation error in terms of the number of servers and stragglers. Comparative analysis demonstrates that our framework significantly outperforms Berrut-based methods for various non-polynomial functions.
I Introduction
Distributed computing is an effective paradigm for executing large-scale computational tasks across multiple parallel computing machines. In such systems, the data sets is distributed among several worker nodes, while a master node coordinates the computation and aggregates the results [1]. A major performance bottleneck in distributed learning systems is the presence of stragglers, i.e., slow or unresponsive worker nodes, which can significantly delay the overall computation. To address this challenge, coded computing has been proposed as a fault-tolerant and efficient approach that mitigates the impact of stragglers by introducing structured redundancy into distributed computations. In particular, by distributing encoded shares of the data among worker nodes, coded computing enables the master node to recover the desired result from the outputs of only a subset of workers, thereby avoiding the need to wait for stragglers. This idea has been successfully applied to problems such as distributed matrix multiplication [2, 3] and polynomial function computation [4, 5, 6]. However, many existing coded computing schemes such as Lagrange Coded Computing(LCC) [4], Analog Lagrange Coded Computing(ALCC) [6] are limited to polynomial computations and rely on exact interpolation, which can be numerically unstable over real-valued data [7] and often require a large number of worker nodes. In contrast, [8] proposed a numerically stable coded computing framework based on Chebyshev polynomials; however, this approach remains restricted to polynomial computations and due to its reliance on exact interpolation, still requires a large number of workers as the degree of polynomial increases.
To address these challenges, Berrut Approximated Coded Computing (BACC) [9] was recently proposed as a numerically stable framework for approximately computing arbitrary (possibly non-polynomial) functions in a distributed setting. BACC employs Berrut rational function for both data encoding and decoding, thereby extending existing coded computing frameworks beyond polynomial functions while providing provable approximation error guarantees. Although BACC substantially improves numerical stability, its decoding procedure relies on global approximation, which limits the achievable approximation order as the number of stragglers increases. These limitations motivate the need to investigate new coded computing designs that improve numerical accuracy while explicitly characterizing the associated trade-offs with respect to computation, communication, and straggler tolerance. In parallel, Basis splines have been extensively studied in approximation theory and numerical analysis as a powerful tool for interpolation and function approximation [10, 11][12][13, 14]. Since the foundational work of Schoenberg, B-splines have been shown to possess several desirable properties, including local support, high-order approximation capability, and strong numerical stability. Despite this rich theoretical foundation, spline-based methods have not been systematically explored in the context of coded computing, where existing approaches continue to rely predominantly on polynomial or rational interpolation schemes.
In this work, we bridge this gap by proposing a B-spline assisted coded computing (BSCC) framework for distributed computation in the presence of stragglers. Unlike global interpolation-based schemes, BSCC employs B-spline basis functions to perform local, piecewise-polynomial approximation of the target function, enabling higher-order approximation accuracy and improved numerical stability. Our specific contributions are listed below:
1) For coded computing frameworks, we formulate a reconstruction methodology based on clamped B-spline interpolation, wherein the composite function evaluations returned by the workers are represented using B-spline basis functions. Our methodology explicitly specifies the construction of the spline space, the use of clamped knot vectors, and the incorporation of natural boundary conditions, resulting in a well-posed linear system for reconstruction at the master node. We analyze the structure of the resulting B-spline interpolation system, and demonstrate that the reconstruction can be carried out efficiently using banded linear solvers, with computational complexity that scales linearly with the number of workers.
2) We establish explicit bounds on the Lebesgue constant associated with the proposed spline-based decoding scheme and connect these bounds to the overall approximation error of the distributed computation. This analysis yields a precise characterization of how the reconstruction error depends on the number of straggling workers.
3) Through both theoretical comparison and numerical simulations, we show that the proposed B-spline–based coded computing framework achieves strictly better approximation accuracy than Berrut Approximated Coded Computing (BACC), exhibiting significantly faster error decay as the number of stragglers increases (see Section IV).
II Preliminaries on B-Splines
In this subsection, we revisit some basic definitions and fundamental properties of B-spline functions.
Definition 1 (Clamped (open) knot vector[11]).
Let $a,b]\subset\mathbb{R}$ be a compact interval, and let $\{\alpha_{i}\}_{i=0}^{N-1}\subset[a,b]$ be a strictly increasing sequence of evaluation points, i.e., $a=\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N-1}=b$. For a given spline degree $p\geq 0$, spline interpolation of $N$ data points requires a non-decreasing knot sequence of length $N+2p$. A knot sequence is said to be $(p+1)$-open (or clamped) on $[a,b]$ if it is non-decreasing and has end knots of multiplicity $p+1$, i.e., $a=t_{0}=\cdots=t_{p}<t_{p+1}\leq\cdots\leq t_{N+p-2}<t_{N+p-1}=\cdots=t_{N+2p-1}=b$. Throughout this work, we employ the $(p+1)$-open (clamped) knot vector $t=(t_{j})_{j=0}^{N+2p-1}$ defined by
$t=[\underbrace{\alpha_{0},\ldots,\alpha_{0}}_{p+1},\alpha_{1},\alpha_{2},\ldots,\alpha_{N-2},\underbrace{\alpha_{N-1},\ldots,\alpha_{N-1}}_{p+1}].$ (1)
Definition 2 (B-spline basis [10]).
Given the knot vector $t$ and spline degree $p\geq 0$, the B-spline basis functions $\{B_{j,p}\}_{j=0}^{N+p-2}$ associated with $\mathcal{S}_{p,t}$ are defined recursively by the Cox–de Boor formula. For $p=0$,
$B_{j,0}(z)=\begin{cases}1,&z\in[t_{j},t_{j+1}),\\ 0,&\text{otherwise},\end{cases}$
and for $p\geq 1$,
$B_{j,p}(z)=\frac{z-t_{j}}{t_{j+p}-t_{j}}B_{j,p-1}(z)+\frac{t_{j+p+1}-z}{t_{j+p+1}-t_{j+1}}B_{j+1,p-1}(z).$
Definition 3 (Support and smoothness of B-splines[12]).
Each B-spline basis function $B_{j,p}$ belongs to $C^{p-1}(a,b)$ and has compact support $\operatorname{supp}(B_{j,p})=(t_{j},t_{j+p+1})$. Moreover, $B_{j,p}(z)\geq 0$ for all $z\in(t_{j},t_{j+p+1})$. Consequently, each basis function is nonzero only over $p+1$ consecutive knot intervals.
Definition 4 (Spline space).
Let $t$ be the clamped knot vector defined in Definition 1. The spline space $\mathcal{S}_{p,t}$ consists of all piecewise polynomial functions of degree at most $p$ defined on $[a,b]$ with smoothness determined by the multiplicities of the knots.
Remark 1 (Effect of knot multiplicity[11]).
For a spline of degree $p$, a knot of multiplicity $m$ induces $C^{p-m}$ smoothness at that knot. In particular, the boundary knots in the clamped knot vector have multiplicity $p+1$, resulting in $C^{-1}$ regularity at the endpoints, while interior knots of unit multiplicity ensure $C^{p-1}$ smoothness.
Definition 5 (B-spline representation of spline interpolant [12]).
Let $f\in C[a,b]$ be a given function, and let $\{\alpha_{i}\}_{i=0}^{N-1}\subset[a,b]$ be a strictly increasing set of evaluation points. Let $t$ be a prescribed knot vector and let $\mathcal{S}_{p,t}:=\operatorname{span}\{B_{0,p},B_{1,p},\ldots,B_{N+p-2,p}\}$ denote the associated spline space of degree $p$, where $\{B_{j,p}\}_{j=0}^{N+p-2}$ are the corresponding B-spline basis functions. A spline interpolant $s\in\mathcal{S}_{p,t}$ is expressed as
$s(z)=\sum_{j=0}^{N+p-2}C_{j,p}\,B_{j,p}(z),$ (2)
where the coefficients $\{C_{j,p}\}$ are determined by enforcing the interpolation conditions $s(\alpha_{i})=f(\alpha_{i}),\quad i=0,\ldots,N-1.$ Substituting (2) into the interpolation conditions yields the linear system
$\sum_{j=0}^{N+p-2}C_{j,p}\,B_{j,p}(\alpha_{i})=f(\alpha_{i}),\qquad i=0,\ldots,N-1.$ (3)
III B-Spline Assisted Coded Computing
We propose a B-spline assisted coded computing (BSCC) framework for approximate computation of an arbitrary function $f(\cdot)$ over a given input dataset. Our setup consists of a master node and $N$ worker nodes indexed by the set $\mathcal{W}=\{\mathcal{W}_{0},\mathcal{W}_{1},\ldots,\mathcal{W}_{N-1}\}$. We assume that each worker node communicates exclusively with the master node through a dedicated communication link, and no communication is permitted among workers. The input dataset is stored at the master node, while all workers are assumed to have prior knowledge of the function to be evaluated. Among the worker nodes, we assume the presence of $s$ number of straggling workers, for $1\leq s<N-2$, which may fail to return their computation results within the prescribed time. Let $\mathbf{X}=(\mathbf{X}_{0},\ldots,\mathbf{X}_{K-1})$ denote the dataset held by the master node, where $\mathbf{X}_{j}\in\mathbb{R}^{m\times n}$ for all $j\in[K]$, with $K]\triangleq\{0,1,2,\ldots,K-1\}$. The primary objective of BSCC framework is to approximately evaluate an arbitrary function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ on the matrices in $\mathbf{X}$ in a decentralized manner while guaranteeing tolerable approximation errors against $s$ number of stragglers. In particular, since the function $f(\cdot)$ may be non-polynomial, BSCC aims to enable numerically stable distributed computation of $f(\cdot)$ on $\mathbf{X}_{j}$ for all $j\in[K]$ with bounded approximation error; that is, if $\mathbf{Y}_{j}$ denotes the output reconstructed from the distributed computation corresponding to $\mathbf{X}_{j}$, then $\mathbf{Y}_{j}\approx f(\mathbf{X}_{j})$ for all $j\in[K]$ within a tolerable accuracy loss. Within this setting, we employ a generalized encoding method available in the literature and propose a B-spline assisted numerically stable reconstruction procedure at the master. [^1] Coded computing under additional constraints such as data privacy and Byzantine adversaries are well studied. In this work, we restrict our attention to resilience against straggling workers. Privacy guarantees and Byzantine robustness are straightforward to generalize using existing techniques.
III-A Encoding
We describe the encoding method, specifically highlighting how the dataset $\mathcal{X}$ is mapped into distributed shares for the worker nodes. To encode the dataset $\mathcal{X}$, the master node may employ any suitable interpolatory basis functions $\{\Phi_{j}(\cdot)\}$, such as the Lagrange basis [6] or the Berrut basis functions [9]. These basis functions satisfy the cardinality property, namely $\Phi_{j}(\beta_{i})=1$ for $j=i$ and $\Phi_{j}(\beta_{i})=0$ for $j\neq i$, for the underlying $K$ ordered non-decreasing encoding points $\{\beta_{i}\in\mathbb{R}~|~i=0,1,\ldots,K-1\}$. Using this property, the dataset $\mathcal{X}$ is encoded as a function of an evaluation variable $z$ as
$u(z)=\sum_{j=0}^{K-1}\mathbf{X}_{j}\Phi_{j}(z),$ (4)
where $\{\mathbf{X}_{j}\}_{j=0}^{K-1}\subset\mathcal{X}$ denote the data blocks and $\{\Phi_{j}(\cdot)\}$ are the chosen basis functions.
III-B Distribution of Shares among the Worker Nodes
Once $u(z)$ is constructed as in (4), the master selects a set of $N$ distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}\subset\mathbb{R}$, such that $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N-1}$. For instance, these points could be chosen as Chebyshev nodes of the first kind or second kind. Using these evaluation points, the master node generates the encoded shares by evaluating the encoding function $u(z)$ at each $\alpha_{i}$, resulting in the set of function values $\{\mathbf{Y}_{i}\}_{i=0}^{N-1}$, where $\mathbf{Y}_{i}=u(\alpha_{i}),\quad i\in[N-1].$ Finally, the master distributes the encoded shares to the worker nodes by assigning the share $\mathbf{Y}_{i}$ to worker $\mathcal{W}_{i}$, for $i\in[N-1]$.
III-C Computation at the Worker Nodes
Once the worker $\mathcal{W}_{i}$ receives its evaluation $\mathbf{Y}_{i}$, it computes $f(\mathbf{Y}_{i})=f(u(\alpha_{i}))$, and returns it to the master. In contrast, the straggling workers maynot return these values within the prescribed deadline.
III-D Function Reconstruction using Spline
In the absence of stragglers (i.e., when $s=0$), the master node receives $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$ from $N$ worker nodes, corresponding to a set of distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$. The objective is to reconstruct an approximation of the composite target function $f(u(z))$ over the interval spanned by these points. Since the function $f(\cdot)$ is arbitrary and may be non-polynomial, exact interpolation [6] is generally not feasible or numerically stable. Consequently, we approximate $f(u(z))$ using a basis-spline function $r_{B-spline}(z)\approx f(u(z))$, where
$r_{B-spline}(z)=\sum_{j}C_{j,p}\,B_{j,p}(z),$ (5)
where $\{B_{j,p}(z)\}$ are basis functions of B-spline, and the coefficients $\{C_{j,p}\}$ are obtained using an appropriate interpolation method using the $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$. Using (5), the desired function $f(\mathbf{X}_{j})$ are reconstructed by evaluating the spline at the points $\{\beta_{j}\}_{j=0}^{K-1}$ as $\mathbf{Y}_{j}=r_{B-spline}(\beta_{j})$. Note that these are approximate evaluations, i..e, $\mathbf{Y}_{j}\approx f(\mathbf{X}_{j})$, for $j\in[K-1]$, and the accuracy of this reconstruction depends on the reconstruction accuracy of $r_{B-spline}(z)$ in (5). In the next section, we present our interpolation method to choose the set $\{C_{j,p}\}$.
III-E Clamped B-Spline Interpolation for BSCC
In our framework, we employ cubic B-splines, i.e., $p=3$, to approximate $f(u(z))$. Let $\{B_{j,3}(z)\}_{j=0}^{N+1}$ denote the set of cubic B-spline basis functions, where the spline space $\mathcal{S}_{3,t}\triangleq\operatorname{span}\{B_{0,3}(z),B_{1,3}(z),\ldots,B_{N+1,3}(z)\},$ is constructed over the knot vector $t$ given by
$t=\bigl[\underbrace{\alpha_{0},\ldots,\alpha_{0}}_{4},\alpha_{1},\alpha_{2},\ldots,\alpha_{N-2},\underbrace{\alpha_{N-1},\ldots,\alpha_{N-1}}_{4}\bigr],$ (6)
where $\{\alpha_{i}\}_{i=0}^{N-1}$ are the avaulation points for BSCC. Note that the above knot vector corresponds to a clamped (open) knot vector for cubic B-splines. In general, for splines of degree $p$, the extreme knot points are repeated $p$ times. As a result, in our case, $\alpha_{0}$ and $\alpha_{N-1}$ are replicated thrice. Given $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$ obtained from the worker nodes at a set of distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$, we approximate the composite function $f(u(z))$ by a cubic spline $s(z)\in\mathcal{S}_{3,t}$ of the form
$s(z)=\sum_{j=0}^{N+1}C_{j,3}\,B_{j,3}(z),$ (7)
where $\{C_{j,3}\}_{j=0}^{N+1}$ are the spline coefficients that need to be computed. Towards that direction, the interpolation conditions $s(\alpha_{i})=f(u(\alpha_{i})),\quad i=0,\ldots,N-1,$ lead to a system of linear equations given by $\mathbf{B}_{rect}\mathbf{c}=\mathbf{f},$ where the rectangular B-spline matrix $\mathbf{B}_{\mathrm{rect}}\in\mathbb{R}^{N\times(N+2)}$ given by
$\mathbf{B}_{rect}=\begin{bmatrix}B_{0,3}(\alpha_{0})&B_{1,3}(\alpha_{0})&\cdots&B_{N+1,3}(\alpha_{0})\\ B_{0,3}(\alpha_{1})&B_{1,3}(\alpha_{1})&\cdots&B_{N+1,3}(\alpha_{1})\\ \vdots&\vdots&\ddots&\vdots\\ B_{0,3}(\alpha_{N-1})&B_{1,3}(\alpha_{N-1})&\cdots&B_{N+1,3}(\alpha_{N-1})\\ \end{bmatrix}$ (8)
has full row-rank $N$, and the vectors $\mathbf{c}$ and $\mathbf{f}$ are given by
$\mathbf{c}=\begin{bmatrix}C_{0,3}\\ C_{1,3}\\ \vdots\\ C_{N+1,3}\end{bmatrix},\mathbf{f}=\begin{bmatrix}f(u(\alpha_{0}))\\ f(u(\alpha_{1}))\\ \vdots\\ f(u(\alpha_{N-1}))\\ \end{bmatrix}$. (9)
Consequently, the resulting linear system is underdetermined and admits infinitely many solutions. To obtain a unique solution, p1=2p-1=2 additional constraints are required. Therefore, we impose the natural spline boundary conditions $s^{\prime\prime}(\alpha_{0})=0\quad\text{and}\quad s^{\prime\prime}(\alpha_{N-1})=0,$ where $s^{\prime\prime}(z)$ denotes the double derivative of $s(z)$. Since $s^{\prime\prime}(z)=\sum_{j=0}^{N+1}C_{j,3}\,B^{\prime\prime}_{j,3}(z),$ these boundary conditions yield two additional linear constraints on the spline coefficients [14]. Evaluating at the boundary points gives
$\sum_{j=0}^{N+1}C_{j,3}B^{\prime\prime}_{j,3}(\alpha_{0})=0\quad\text{and}\quad\sum_{j=0}^{N+1}C_{j,3}B^{\prime\prime}_{j,3}(\alpha_{N-1})=0.$
By augmenting the rectangular system with these two boundary constraints, we obtain a square B-spline matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$. The resulting linear system can be written as $\mathbf{B}\mathbf{c}=\bar{\mathbf{f}}$, where $\bar{\mathbf{f}}=[0~\mathbf{f}^{T}~0]^{T}$ and
$\mathbf{B}=\begin{bmatrix}B^{\prime\prime}_{0,3}(\alpha_{0})&B^{\prime\prime}_{1,3}(\alpha_{0})&\cdots&B^{\prime\prime}_{N+1,3}(\alpha_{0})\\ B_{0,3}(\alpha_{0})&B_{1,3}(\alpha_{0})&\cdots&B_{N+1,3}(\alpha_{0})\\ B_{0,3}(\alpha_{1})&B_{1,3}(\alpha_{1})&\cdots&B_{N+1,3}(\alpha_{1})\\ \vdots&\vdots&\ddots&\vdots\\ B_{0,3}(\alpha_{N-1})&B_{1,3}(\alpha_{N-1})&\cdots&B_{N+1,3}(\alpha_{N-1})\\ B^{\prime\prime}_{0,3}(\alpha_{N-1})&B^{\prime\prime}_{1,3}(\alpha_{N-1})&\cdots&B^{\prime\prime}_{N+1,3}(\alpha_{N-1})\end{bmatrix}$. (10)
The two additional zero entries in the right-hand side vector correspond to the imposed natural spline boundary conditions. Before solving this linear system, we next present results on the structure of $\mathbf{B}$ and establish results on its invertibility as well as the computational complexity of its inversion.
Proposition 1.
For cubic B-spline i.e., $p=3$, the B-spline matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$ in (10) is banded and full rank, and hence the associated linear system $\mathbf{B}\mathbf{c}=\bar{\mathbf{f}}$ admits a unique solution.
Proposition 2.
For cubic B-splines, the matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$ in (10) has upper and lower bandwidth $2$. Subsequently, the associated linear system can be solved using banded LU decomposition with computational complexity $\mathcal{O}(8(N+2))$, which is smaller than that of a dense matrix.
Overall, after obtaining the spline coefficients $\{C_{j,3}\}_{j=0}^{N+1}$, the desired function values are reconstructed as explained in Section III.
IV Experimental Results
[…]
Figure 1: Average relative error (in dB) of the BSCC and BACC decoding schemes with Lagrange and Berrut encoding for $f(\mathbf{X})=\mathbf{X}\sin(\mathbf{X})$ and $f(\mathbf{X})=\frac{1}{1+e^{-\mathbf{X}}}$, with parameters $N=100$, $K=8$, and $\mathbf{X}\in\mathbb{R}^{40\times 5}$.
The plots in Fig. 1 confirm that BSCC achieves a remarkable improvement in approximation accuracy over BACC for non-polynomial functions, particularly when the number of stragglers is small, and exhibits a significant improvement even when the number of stragglers increases. Moreover, the results highlight that, to maintain a relative error on the order of $10^{-2}$ or $10^{-3}$, the BSCC framework can tolerate a substantially larger number of stragglers compared to the BACC framework.