### v2

# Basis-Spline Assisted Coded Computing: Strategies and Error Bounds

Abstract
Coded computing has emerged as a key framework for addressing the impact of stragglers in distributed computation. While polynomial functions often admit exact recovery under existing coded computing schemes, non-polynomial functions require approximate reconstruction from a finite number of evaluations, posing significant challenges. Consequently, interpolation-based methods for non-polynomial coded computing have gained attention, with Berrut approximated coded computing emerging as a state-of-the-art approach. However, due to the global support of Berrut interpolants, the reconstruction accuracy degrades significantly as the number of stragglers increases. To address this challenge, we propose a coded computing framework based on cubic B-spline interpolation. In our approach, server-side function evaluations are reconstructed at the master using B-splines, exploiting their local support and smoothness properties to enhance stability and accuracy. We provide a systematic methodology for integrating B-spline interpolation into coded computing and derive theoretical bounds on approximation error for certain class of smooth functions. Our analysis demonstrates that the error bounds of our approach exhibit a faster decay with respect to the number of workers compared to the Berrut-based method. Experimental results also confirm that our method offers improved accuracy over Berrut-based methods for various smooth non-polynomial functions.
I Introduction
Distributed computing is an effective paradigm for executing large-scale computational tasks, wherein a master coordinates the computation by distributing the dataset among several workers [1]. A major performance bottleneck in such systems is the presence of stragglers, i.e., slow or unresponsive workers, which can significantly delay the overall computation. To address this challenge, coded computing has been proposed as a fault-tolerant approach by introducing structured redundancy into distributed computations. This idea has been successfully applied to problems such as distributed matrix multiplication [2, 3] and polynomial function computation [4, 5, 6]. However, many existing coded computing schemes such as Lagrange Coded Computing [4], Analog Lagrange Coded Computing [6] are limited to polynomial computations and rely on exact interpolation, which can be numerically unstable over real-valued data [7] and often require a large number of workers. In contrast, [8] proposed a numerically stable coded computing framework based on Chebyshev polynomials, however, this approach remains restricted to polynomial computations and due to its reliance on exact interpolation, still requires a large number of workers as the degree of polynomial increases.
To address these challenges, Berrut Approximated Coded Computing (BACC) [9] was proposed as a numerically stable framework for approximately computing arbitrary functions. BACC employs Berrut rational functions for both data encoding and reconstruction, thereby extending existing coded computing frameworks beyond polynomial functions while providing provable approximation error guarantees. Although BACC improves numerical stability, its reconstruction procedure relies on global approximation, which limits the achievable approximation order as the number of stragglers increases. To alleviate this problem, [10] explored a learning-theoretic framework to coded computing, wherein the underlying encoding and reconstruction operations were posed as regression problems, and were subsequently solved using data-centric methods. Interestingly, [10] was shown to outperform [9] in terms of relative error when handling non-polynomial functions. Although [10] is the state-of-the art in terms of accuracy metric, it is accompanied by other implementation overheads such as dataset management and model-training. Thus, within the class of non data-centric methods, BACC is still the state-of-the-art, and therefore, there is a need to investigate new structured coded computing methods that outperform numerical accuracy of BACC while explicitly characterizing the associated trade-offs.
I-A Contributions
Basis splines (B-splines) have been extensively studied in approximation theory and numerical analysis as a powerful tool for interpolation and function approximation [11, 12, 13, 14, 15]. In particular, they possess several desirable properties, including local support, higher-order approximation capability, and strong numerical stability. Despite this rich theoretical foundation, spline-based methods have not been systematically explored in the context of coded computing, where existing approaches continue to rely predominantly on polynomial or rational interpolation schemes. To bridge this gap, in this work, we propose a B-spline assisted coded computing (BSCC) framework for distributed computation of arbitrary functions in the presence of stragglers. Unlike global interpolation-based schemes, BSCC employs B-spline basis functions to perform local, piecewise-polynomial approximation of the target function, enabling higher-order approximation accuracy and improved numerical stability.
Our specific contributions are listed below: For coded computing frameworks, we formulate a reconstruction methodology based on cubic B-spline interpolation. Our methodology explicitly specifies the construction of the spline space, the use of clamped knot vectors, and the natural boundary conditions, resulting in a well-posed linear system for reconstruction. We analyze the structure of the resulting B-spline interpolation system, and demonstrate that the reconstruction can be carried out efficiently using banded linear solvers, with computational complexity that scales linearly with the number of workers (see Section III). We establish explicit bounds on the operator norm associated with the proposed B-spline reconstruction and connect these bounds to the overall approximation error. This analysis yields a precise characterization of how the reconstruction error depends on the number of workers and stragglers. In particular, for a class of smooth non-polynomial functions, we show that the approximation error of BSCC exhibits an inverse-quadratic decay with respect to the number of workers (see Section IV). Finally, through experimental results, we show that BSCC outperforms BACC in terms of approximation accuracy for a broader class of smooth non-polynomial functions, however, with a marginal increase in the computational complexity during the reconstruction process (see Section V).
II Preliminaries on B-Splines
In this section, we revisit some basic definitions and fundamental properties of B-spline functions used in the our construction and analysis. To define B-spline functions, we define knot vectors, which capture the support set of B-splines.
Definition 1.
A knot vector is a finite non-decreasing sequence of real numbers, denoted by $\mathbf{t}=\{t_{i}\}_{i=0}^{m-1}$, such that $t_{0}\leq t_{1}\leq\cdots\leq t_{m-1}$, where $m\in\mathbb{N}$. The elements of $\mathbf{t}$ are called knots.
Definition 2.
Let $p\geq 0$, and let $\mathbf{t}=\{t_{j}\}_{j=0}^{m-1}$ be a knot vector. Provided that $m\geq p+2$, the B-spline basis functions of degree $p$, denoted by $\{B_{j,p}(z)\mid j=0,1,\ldots,m-p-2\}$, are defined recursively over the knot vector $\mathbf{t}$ [11] as follows:
For $p=0$,
$B_{j,0}(z)=\begin{cases}1,&z\in[t_{j},t_{j+1}),\\ 0,&\text{otherwise},\end{cases}$
and for $p\geq 1$,
$B_{j,p}(z)=\frac{z-t_{j}}{t_{j+p}-t_{j}}B_{j,p-1}(z)+\frac{t_{j+p+1}-z}{t_{j+p+1}-t_{j+1}}B_{j+1,p-1}(z).$
If a denominator is zero, the corresponding term is defined to be zero.
Each B-spline basis function $B_{j,p}(z)$ satisfies $B_{j,p}(z)\geq 0$ for all $z$ and $B_{j,p}(z)=0$ for $z\notin[t_{j},t_{j+p+1})$. Consequently, each $B_{j,p}(z)$ has compact support and is nonzero only on the union of at most $p+1$ consecutive knot intervals. In the context of interpolation and approximation, B-spline basis functions are also defined over a special class of knot vectors called clamped knot vectors [12], which are formally defined below.
Definition 3.
Let $a,b]\subset\mathbb{R}$ be a compact interval and let $p\geq 0$ and $n>p$ be integers. Let $a=t_{0}<t_{1}<\cdots<t_{n-1}=b$ be a strictly increasing sequence of real numbers. A knot vector $\mathbf{t}$ is said to be $(p+1)$-clamped on $[a,b]$ if it is defined as
$\mathbf{t}=[\underbrace{t_{0},\ldots,t_{0}}_{p+1},t_{1},\ldots,t_{n-2},\underbrace{t_{n-1},\ldots,t_{n-1}}_{p+1}].$
Definition 4.
For $p\!\geq\!0$ and $m\!\geq\!p\!+\!2$, let $\{B_{j,p}(z)\mid j=0,1,\ldots,m\!-\!p\!-\!2\}$ denote the B-spline basis functions of degree $p$ defined with respect to a knot vector $\mathbf{t}=\{t_{i}\}_{i=0}^{m-1}$. Using these basis functions a spline space denoted by $\mathcal{S}_{p,\mathbf{t}}$, can be generated as[13]
$\mathcal{S}_{p,\mathbf{t}}=\bigg\{\sum_{j=0}^{m-p-2}C_{j,p}\,B_{j,p}(z)\;\forall\;C_{j,p}\in\mathbb{R}\bigg\}.$ (1)
The B-spline basis functions $\{B_{j,p}(z)\mid j=0,1,\ldots,m\!-\!p\!-\!2\}$ satisfy the partition of unity property, i.e., $\sum_{j=0}^{m-p-2}B_{j,p}(z)=1$, over its support set.
III B-Spline Assisted Coded Computing
We propose a B-spline assisted coded computing (BSCC) framework for approximate computation of an arbitrary function $f(\cdot)$ over a given input dataset. Our setup consists of a master and $N$ workers indexed by the set $\mathcal{W}=\{\mathcal{W}_{0},\mathcal{W}_{1},\ldots,\mathcal{W}_{N-1}\}$. We assume that each worker communicates exclusively with the master through a dedicated communication link, and no communication is permitted among workers. The input dataset is stored at the master, while all workers are assumed to have prior knowledge of the function to be evaluated. Among the workers, we assume the presence of $S$ number of stragglers, for $1\leq S<N-2$, which may fail to return their computation results within the prescribed time. Let $\mathbf{X}=(\mathbf{X}_{0},\ldots,\mathbf{X}_{K-1})$ denote the dataset held by the master, where $\mathbf{X}_{j}\in\mathbb{R}^{m\times n}$ for all $j\in[K]$, with $K]\triangleq\{0,1,2,\ldots,K-1\}$. The primary objective of BSCC framework is to approximately evaluate an arbitrary function $f:\mathbb{R}^{m\times n}\rightarrow\mathbb{R}^{m\times n}$ on the matrices in $\mathbf{X}$ in a decentralized manner while guaranteeing tolerable approximation errors against $S$ number of stragglers. In particular, since the function $f(\cdot)$ may be non-polynomial, BSCC aims to enable numerically stable distributed computation of $f(\cdot)$ on $\mathbf{X}_{j}$ for all $j\in[K]$ with bounded approximation error; that is, if $\mathbf{Y}_{j}$ denotes the output reconstructed from the distributed computation corresponding to $\mathbf{X}_{j}$, then $\mathbf{Y}_{j}\approx f(\mathbf{X}_{j})$ for all $j\in[K]$ within a tolerable accuracy loss. Within this setting, we employ a generalized encoding method available in the literature and propose a B-spline assisted numerically stable reconstruction procedure at the master. [^1] Coded computing under additional constraints such as data privacy and Byzantine adversaries are well studied. In this work, we restrict our attention to resilience against stragglers. Privacy guarantees and Byzantine robustness are straightforward to generalize using existing techniques.
III-A Encoding
We describe the encoding method, specifically highlighting how the dataset $\mathcal{X}$ is mapped into distributed shares for the workers. To encode the dataset, the master may employ any suitable interpolatory basis functions $\{\Phi_{j}(\cdot)\}$, $j\in[K]$, such as the Lagrange basis [4] or the Berrut basis functions [9]. These basis functions satisfy the cardinality property, namely $\Phi_{j}(\beta_{i})=1$ for $j=i$ and $\Phi_{j}(\beta_{i})=0$ for $j\neq i$, for the underlying $K$ ordered non-decreasing encoding points $\{\beta_{i}\in\mathbb{R}~|~i=0,1,\ldots,K-1\}$. Using this property, the dataset $\mathcal{X}$ is encoded as a function of an evaluation variable $z$ as
$u(z)=\sum_{j=0}^{K-1}\mathbf{X}_{j}\Phi_{j}(z),$ (2)
where $\{\mathbf{X}_{j}\}_{j=0}^{K-1}\subset\mathcal{X}$ denote the data blocks and $\{\Phi_{j}(\cdot)\}$ are the chosen basis functions.
III-B Distribution of Shares among the Workers
Once $u(z)$ is constructed as in (2), the master selects a set of $N$ distinct and ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}\subset\mathbb{R}$, such that $\alpha_{0}<\alpha_{1}<\cdots<\alpha_{N-1}$. For instance, these points could be chosen as Chebyshev nodes of the first kind or second kind. Using these evaluation points, the master generates the encoded shares by evaluating the encoding function $u(z)$ at each $\alpha_{i}$, resulting in the set of function values $\{\mathbf{Y}_{i}\}_{i=0}^{N-1}$, where $\mathbf{Y}_{i}=u(\alpha_{i}),\quad i\in[N].$ Finally, the master distributes the encoded shares to the workers by assigning the share $\mathbf{Y}_{i}$ to worker $\mathcal{W}_{i}$, for $i\in[N]$.
III-C Computation at the Workers
Once the worker $\mathcal{W}_{i}$ receives its evaluation $\mathbf{Y}_{i}$, it computes $f(\mathbf{Y}_{i})=f(u(\alpha_{i}))$, and returns it to the master. In contrast, the stragglers may not return these values within the prescribed deadline.
III-D Function Reconstruction using B-spline
In the absence of stragglers, i.e., when $S=0$, the master receives $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$ from $N$ workers, corresponding to the ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$. The objective of the master is to reconstruct an approximation of the composite function $f(u(z))$ over the interval spanned by these points. Since the function $f(\cdot)$ is arbitrary and may be non-polynomial, exact interpolation [6] is generally not feasible or numerically stable. Consequently, we approximate $f(u(z))$ using a B-spline function $r_{B-spline}(z)\approx f(u(z))$, where
$r_{B-spline}(z)=\sum_{j}C_{j,p}\,B_{j,p}(z),$ (3)
where $\{B_{j,p}(z)\}$ are B-spline basis functions of degree $p$, and the coefficients $\{C_{j,p}\}$ are obtained using an appropriate interpolation method using the $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$. Using (3), the desired function $f(\mathbf{X}_{j})$ are reconstructed by evaluating the spline at the points $\{\beta_{j}\}_{j=0}^{K-1}$ as $\mathbf{Y}_{j}=r_{B-spline}(\beta_{j})$. Note that these are approximate evaluations, and the accuracy of this reconstruction depends on the reconstruction accuracy of $r_{B-spline}(z)$ in (3). In the next section, we present our interpolation method to choose the set $\{C_{j,p}\}$ using cubic B-splines, i.e., p=3p=3.222In approximation theory, it is well known that among various degrees of B-spline, cubic B-splines achieve a good balance between the accuracy of approximation and computational complexity for interpolation.
III-E Cubic B-Spline Interpolation for BSCC
With the use of cubic B-splines, i.e., p=3\!p=\!3, let $\{B_{j,3}(z)\}_{j=0}^{N+1}$ denote the set of cubic B-spline basis functions, where the spline space $\mathcal{S}_{3,\mathbf{t}}$ is constructed over the $4$-clamped knot vector $\mathbf{t}$ given by
$\mathbf{t}=\bigl[\underbrace{\alpha_{0},\ldots,\alpha_{0}}_{4},\alpha_{1},\alpha_{2},\ldots,\alpha_{N-2},\underbrace{\alpha_{N-1},\ldots,\alpha_{N-1}}_{4}\bigr],$ (4)
where $\{\alpha_{i}\}_{i=0}^{N-1}$ are the distinct and ordered evaluation points for BSCC. Given $N$ function evaluations $\{f(u(\alpha_{i}))\mid i=0,1,\ldots,N-1\}$, we approximate the composite function $f(u(z))$ by a cubic spline $s(z)\in\mathcal{S}_{3,\mathbf{t}}$ of the form
$s(z)=\sum_{j=0}^{N+1}C_{j,3}\,B_{j,3}(z),$ (5)
where $\{C_{j,3}\}_{j=0}^{N+1}$ are the spline coefficients that need to be computed. Towards that direction, the interpolation conditions $s(\alpha_{i})=f(u(\alpha_{i})),\quad i=0,\ldots,N-1,$ lead to a system of linear equations given by $\mathbf{B}_{rect}\mathbf{c}=\mathbf{f},$ where the rectangular B-spline matrix $\mathbf{B}_{\mathrm{rect}}\in\mathbb{R}^{N\times(N+2)}$ given by
$\mathbf{B}_{rect}=\begin{bmatrix}B_{0,3}(\alpha_{0})&B_{1,3}(\alpha_{0})&\cdots&B_{N+1,3}(\alpha_{0})\\ B_{0,3}(\alpha_{1})&B_{1,3}(\alpha_{1})&\cdots&B_{N+1,3}(\alpha_{1})\\ \vdots&\vdots&\ddots&\vdots\\ B_{0,3}(\alpha_{N-1})&B_{1,3}(\alpha_{N-1})&\cdots&B_{N+1,3}(\alpha_{N-1})\\ \end{bmatrix}$ (6)
has full row-rank $N$, and the vectors $\mathbf{c}$ and $\mathbf{f}$ are given by
$\mathbf{c}=\begin{bmatrix}C_{0,3}\\ C_{1,3}\\ \vdots\\ C_{N+1,3}\end{bmatrix},\mathbf{f}=\begin{bmatrix}f(u(\alpha_{0}))\\ f(u(\alpha_{1}))\\ \vdots\\ f(u(\alpha_{N-1}))\\ \end{bmatrix}$. (7)
Consequently, the resulting linear system is underdetermined and admits infinitely many solutions. To obtain a unique solution, p1=2p\!-1\!=2 additional constraints are required. Therefore, we impose the natural spline boundary conditions $s^{\prime\prime}(\alpha_{0})=0\quad\text{and}\quad s^{\prime\prime}(\alpha_{N-1})=0,$ where $s^{\prime\prime}(z)$ denotes the double derivative of $s(z)$. Since $s^{\prime\prime}(z)=\sum_{j=0}^{N+1}C_{j,3}\,B^{\prime\prime}_{j,3}(z),$ these boundary conditions yield two additional linear constraints on the spline coefficients [15]. Evaluating them gives
$\sum_{j=0}^{N+1}C_{j,3}B^{\prime\prime}_{j,3}(\alpha_{0})=0\quad\text{and}\quad\sum_{j=0}^{N+1}C_{j,3}B^{\prime\prime}_{j,3}(\alpha_{N-1})=0.$
By augmenting the rectangular system with these two boundary constraints, we obtain an $(N+2)\times(N+2)$ square B-spline matrix. The resulting system of linear equations can be written as $\mathbf{B}\mathbf{c}=\bar{\mathbf{f}}$, where $\bar{\mathbf{f}}=[0~\mathbf{f}^{T}~0]^{T}$ and
$\mathbf{B}=\begin{bmatrix}\mathbf{a}\\ \mathbf{B}_{\mathrm{rect}}\\ \mathbf{b}\end{bmatrix}$, (8)
where the row vectors $\mathbf{a},\mathbf{b}\in\mathbb{R}^{1\times(N+2)}$are given by
$\mathbf{a}=\bigl[B^{\prime\prime}_{0,3}(\alpha_{0}),\;B^{\prime\prime}_{1,3}(\alpha_{0}),\;\ldots,\;B^{\prime\prime}_{N+1,3}(\alpha_{0})\bigr],$
$\mathbf{b}=\bigl[B^{\prime\prime}_{0,3}(\alpha_{N-1}),\;B^{\prime\prime}_{1,3}(\alpha_{N-1}),\;\ldots,\;B^{\prime\prime}_{N+1,3}(\alpha_{N-1})\bigr].$
The two additional zero entries in $\bar{\mathbf{f}}$ correspond to the imposed natural spline boundary conditions. Before solving this linear system, we next present results on the structure of $\mathbf{B}$ and establish results on its invertibility as well as the computational complexity for its inversion.
Proposition 1.
[…]
Proof.
See Appendix -A for the proof. ∎
Proposition 2.
For cubic B-splines, the matrix $\mathbf{B}\in\mathbb{R}^{(N+2)\times(N+2)}$ in (8) has upper and lower bandwidth of $2$. Subsequently, the associated linear system can be solved using banded LU decomposition with computational complexity $\mathcal{O}(8(N+2))$[17], which is smaller than that of a dense matrix.
Overall, after obtaining the spline coefficients $\{C_{j,3}\}_{j=0}^{N+1}$, the desired function values are reconstructed as explained in Section III-D. In the rest of the section, we discuss the modifications required for interpolation in the presence of stragglers, i.e., when $S>0$. Let $\mathcal{F}\!\!=\!\!\{l_{1},l_{2},\ldots,l_{M}\}$ denote the index set of the M=NSM\!\!=\!\!N-S non-stragglers. Let $\{\mathbf{R}_{l_{1}},\mathbf{R}_{l_{2}},\ldots,\mathbf{R}_{l_{M}}\}$ denote the corresponding results received at the master, where $\mathbf{R}_{l_{i}}=f\!\left(u(\alpha_{l_{i}})\right)$ for $l_{i}\in\mathcal{F}$. Let $\{\alpha_{l_{i}}\}_{i=0}^{M-1}\subset\{\alpha_{i}\}_{i=0}^{N-1}$ denote the ordered set of evaluation points corresponding to the non-stragglers in the set $\mathcal{F}$. To generate the B-spline basis functions, the master constructs the associated $4$-clamped knot vector from evaluation points of the non-stragglers set. Subsequently, $\mathbf{B}_{rect}$ defined in (6) reduces to the dimension $M\times(M+2)$, where the basis functions $\{B_{i,3}(z)\}_{i=0}^{M+1}$ are evaluated at the points $\{\alpha_{l_{i}}\}_{i=0}^{M-1}$. Further, after imposing the boundary conditions on $\mathbf{B}_{rect}$, we obtain $\mathbf{B}$ defined in (8) of dimension $(M+2)\times(M+2)$. Finally, solving the resulting linear system yields the spline coefficients, which are subsequently used to recover the desired values $\{f(\mathbf{X}_{j})\}_{j=0}^{K-1}$. Similar to BACC, there is no strict recovery threshold for BSCC as well. The master reconstructs the function using the evaluations from the non-stragglers, with accuracy improving as more evaluations are received.
IV Error Bounds for BSCC
As discussed in Section III-D, the reconstruction accuracy of $\{f(\mathbf{X}_{j})\}_{j=0}^{K-1}$ depends on the accuracy with which $r_{B-spline}(z)$ approximates $f(u(z))$. Therefore, we are interested in deriving bounds on the approximation error $r_{B-spline}(z)-f(u(z))||_{\infty}$, where $||\cdot||_{\infty}$ denotes the infinity norm over the support set. Henceforth, we use $g(z)$ to refer to $f(u(z))$ and $Tg(z)$ to refer to $r_{B-spline}(z)$, assuming that $T$ is the underlying B-spline interpolation operator that approximates $g(z)$ to $Tg(z)\in\mathcal{S}_{3,\mathbf{t}}$. Towards evaluating the interpolator operator $T$, we recollect this definition from [11].
Definition 5.
Let $C[a,b]$ represent the set of all continuous functions defined over $[a,b]$. Further, for any function $g(z)\in C[a,b]$, let $Tg(z)\in\mathcal{S}_{3,\mathbf{t}}$ denote the spline function obtained by applying the operator $T$ to $g(z)$. Then the operator norm of $T$ is defined as
$\|T\|:=\sup_{g(z)\neq 0}\frac{\|Tg(z)\|_{\infty}}{\|g(z)\|_{\infty}}.$ (9)
For the proposed BSCC interpolation with cubic B-splines, we derive an upper bound on the operator norm $T||$, which is presented in the following proposition.
Proposition 3.
For a given $N$, consider the BSCC setting when $S=0$. Let $T$ denote our spline interpolation operator that maps the composite function $g(z)$ to a spline function $Tg(z)\in\mathcal{S}_{3,\mathbf{t}}$. Then, $\|T\|$ is upper bounded as
$\|T\|\leq\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty},$ (10)
where $\mathbf{B}^{-1}]_{\mathrm{sub}}\!\in\!\mathbb{R}^{(N+2)\times{N}}$ denotes the submatrix of $\mathbf{B}^{-1}$ after excluding the first and the last columns, where $\mathbf{B}$ is given in (8).
Proof.
See Appendix -B for the proof. ∎
For a given set of distinct ordered evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$, the above upper bound $\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}$ can be computed using (8) in a straightforward manner. However, in the following theorem, we derive an upper bound on $\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}$ only as a function of the knot spacings.
Theorem 1.
For a given $N$, a knot vector $\mathbf{t}$, when S=0S\!\!\!=\!\!\!0, let $h_{\max}\!\!\!=\!\!\!\max_{\begin{subarray}{c}0\leq i\leq N-2\\ \end{subarray}}\left(t_{i+1}-t_{i}\right)$ and $h_{\min}\!\!=\!\!\min_{\begin{subarray}{c}0\leq i\leq N-2\end{subarray}}\left(t_{i+1}-t_{i}\right)$ represent the maximum and minimum knot spacing between consecutive distinct knots in the knot vector. Then
$\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}\leq C_{1}(N+2)\frac{h_{\max}}{h_{\min}},$
where $C_{1}>0$ is a constant, which depends only on the degree of the spline.
Proof.
See Appendix -C for the proof. ∎
Further, using the above upper bound on $\|[\mathbf{B}^{-1}]_{\mathrm{sub}}\|_{\infty}$, we use the results from [11] to present an upper bound on the approximation error of BSCC.
Theorem 2.
For a given $N$, consider the BSCC scheme when $S=0$. Then, for every function $g(z)\in C[a,b]$, the interpolation error of BSCC satisfies
$\|g(z)-Tg(z)\|_{\infty}$ $\leq\Bigl(1+C_{1}(N+2)\frac{h_{\max}}{h_{\min}}\Bigr)D_{S_{3,t}}(g(z)),$ (11)
where $D_{S_{3,t}}(g(z))=\inf_{s(z)\in S_{3,\mathbf{t}}}\|g(z)-s(z)\|_{\infty}$, and all other constants are as defined in Theorem 1.
Proof.
See Appendix -D for the proof. ∎
While the above theorem gives a bound on the approximation error of BSCC for an arbitrary function $g(z)$, there is no simpler representation of $D_{S_{3,t}}(z)$ in general. However, for a special class of functions, i.e., when $g(z)\in C^{4}[a,b]$, where $C^{4}[a,b]$ denotes the set of all functions over $[a,b]$ whose fourth derivative is also continuous, we use the results in [11, Chapter XII, page 149] to bound $D_{S_{3,t}}(g(z))$, and subsequently derive an upper bound on (11).
Corollary 1.
For the BSCC setting considered in Theorem 2, if $g(z)\in C^{4}[a,b]$, then there exists a constant $C>0$, depending only on the degree of spline, such that
$\|g(z)-Tg(z)\|_{\infty}\leq C\bigg(1+C_{1}(N+2)\frac{h_{\max}}{h_{\min}}\bigg)\,h_{max}^{4}\|g^{\prime\prime\prime\prime}(z)\|_{\infty},$ (12)
where $g^{\prime\prime\prime\prime}(z)$ represents the fourth derivative of $g(z)$ in $[a,b]$, and all other constants are as defined in Theorem 2.
All the above results are presented when $S=0$. However, with $S>0$, Theorem 2 and Corollary 1 continue to hold with the exception that $N$ must be replaced by NSN-S and the parameters $h_{max}$ and $h_{min}$ must be computed using the evaluation indices of non-stragglers.
IV-A Error Bounds Comparison with BACC
Although BSCC allows any choice of evaluation points $\{\alpha_{i}\}_{i=0}^{N-1}$ over $[a,b]$, one can select $\{\alpha_{i}\}_{i=0}^{N\!-\!1}$ as Chebyshev nodes of the second kind, defined by $\alpha_{i}=\cos\!\left(\frac{i\pi}{N}\right)$, for $i\in[N]$. In such a case, with $S$ stragglers, the upper bound in (12) reduces to
$2C\left(\sin^{4}\!\left(\frac{(S+1)\pi}{2N}\right)+\frac{C_{1}(N-S+2)}{h_{min}}\sin^{5}\!\left(\frac{(S+1)\pi}{2N}\right)\right)\|g^{\prime\prime\prime\prime}(z)\|_{\infty},$ (13)
where $h_{min}$ is independent of $S$, and is known to follow $\mathcal{O}(\frac{1}{N^{2}})$ behavior, for the class of Chebyshev nodes. We recall from [9, Theorem 9] that the approximation error of BACC varies with $S$ and $N$ as
$\bigg(1+\frac{(1+S)(3+S)\pi^{2}}{4}\bigg)\sin\!\left(\frac{(S+1)\pi}{2N}\right),$ (14)
Figure 1: Average relative error (in dB) of the BSCC and BACC schemes with Lagrange and Berrut encoding with N=100N\!=\!100, $K=8$.
after excluding all other constants. Provided $g(z)\in C^{4}[a,b]$, using (13) and (14), we make the following inferences: With $S=0$, the error bound of BACC varies as $\mathcal{O}(\frac{1}{N})$, whereas the error bound of BSCC varies as $\mathcal{O}(\frac{1}{N^{2}})$, thereby indicating faster decay of approximation error with the number of workers. For $0<S<<N$, in the regime of large $N$, the error bound of BACC behaves as $\mathcal{O}(\frac{S^{3}}{N})$, whereas the error bound of BSCC varies as $\mathcal{O}(\frac{S^{5}}{N^{2}})$. This indicates that, for small $S$ and large $N$, the absolute error values of BSCC are expected to be much lower than that of BACC owing to the inverse-quadratic behavior with respect to $N$. Furthermore, along the lines of BACC, for a fixed $N$, we observed that the error bound of BSCC also grows with respect to $S$. In the next section, we attempt to validate these observations using experimental results for arbitrary functions.
V Experimental Results
[…]
VI Future work
In this work, we have proposed a B-spline-assisted coded computing framework applicable to distributed computation of arbitrary functions. Our future work is to extend the proposed framework to incorporate Byzantine workers, develop robust encoding strategies based on B-splines, and adopt the proposed method to accelerate distributed computational tasks in a wide range of applications, including distributed learning.